the greatest possible number of points of intersection of 8 straight lines and 4 circles is

Grade:12th pass

1 Answers

Khimraj
6 years ago
There will be 3 type of intersection point.
  1. between a line and another line = 8C2*1 = 28.
  2. between a circle and another circle =  4C2*2 = 12.
  3. between a circle and a line = 8C1*4C1*2 = 64.
So maximum no of intersection points are 28+12+64 = 104.
Hope it clears. If you like answer then please approve it.